The sum of 3 numbers in an Arithmetic sequence is 12 and the sum of their cubes is 288. Find the numbers.

Grade:10

1 Answers

Rajat
4 years ago
Let the numbers be a-d, a, a+d
So, 3a= 12 
Or, a= 4
(4-d)^3+64+(4+d)^3=288
Or, 24d^2= 96 
Or, d^2 = 4 
Or, d= ±2
Numbers are
2,4,6
Please approve